A fossil gives 435 disintegrations per g of carbon during 1 h of counting. The age of the fossil given that the living matter contains 1.5x10-12 g of 14C per g of carbon and that the half life of beta decay of 14C is 5680yr, will be


1
Expert's answer
2021-01-07T05:31:04-0500

Nt = Noe-"\lambda" t


"\dfrac{Nt}{No} =" e-"\lambda" t (but Nt/No = 1/2)


"\dfrac{1}{2} =" e-"\lambda" t


to cancel e


ln ("\dfrac{1}{2} ) =" ln(e-"\lambda" t )


ln ("\dfrac{1}{2}" ) = -"\lambda"t


-0.693 = -"\lambda"t



Also

Given ;

Mass of 14C = 1.5x10-12g


No of moles


"Moles = \dfrac{mass}{RAM} = \dfrac{1.5x10^-12g}{14g/mol}"


n = 1.07x10-13moles


Number of atoms of 14C in living matter = n x NA


= 1.07x10-13moles x 6.022x10-23 atoms


N= 6.445x1010 atoms


Activity in living matter is given by;

"(\dfrac{-dN}{dt}) = \lambda"N = "\dfrac{0.693N}{t1/2}"


Since t1/2 = 5680yrs


Activity of 14C in living matter is given by;


"\lambda" "=\dfrac{0.693}{t1/2}N"


"\lambda = \dfrac{0.693}{5680 x 365 x 24hrs} x 6.445x10^10"


= 897.6 disintegrations per gram per hour


Now we have

N0 = initial activity = 897.6 disintegrations per hr

Nt = final activity = 435 disintegrations per hr


From the equation;


t1/2 = "\dfrac{0.693}{\lambda}"


"\lambda" = "\dfrac{0.693}{5680 yrs} = 1.22x10^-4 yrs"


Now, from the equation

Nt = Noe-"\lambda" t


"\dfrac{Nt}{No} =" e-"\lambda" t


"\dfrac{435}{897.6} = e" -(1.22x10-4) x t



0.4846 = e-1.22x10-4 x t


t = ln ("\dfrac{0.4846}{-122x10^-4})"


t = 5937.629 years